A spring, whose unstretched length is 2.0 m is compressed to a length of 1/2 m

when a force of 10 N is applied. Find the work required to compress the spring

to a length of 1.0 m.


1
Expert's answer
2022-05-12T07:29:26-0400

-10=k(0.5-2)=-1.5k

k=6.67

F(x)=6.67 x

"W=\int_0^{-1}6.67x dx=6.67 x^2/2|^{-1}_0=3.34(1-0)=3.34J"